A jetliner, traveling northward, is landing with a speed of 59.5 m/s. Once the jet touches down, it has 787 m of runway in which to reduce its speed to 5.75 m/s. Compute the average acceleration (magnitude and direction) of the plane during landing (take the direction of the plane's motion as positive).
Answer #1

answer) from kinematics we have

vf2=vi2-2as

a=vf2-vi2/-2s

putting the values

a=59.52-5.752/-2*787=2.23 m/s2

so the answer is 2.23 m/s2 and direction is south.

or you can write it like this

a=2.23m/s2 -j